Given the following demand curves facing a monopoly : Q = 14 - 2P Find the profit maximising price and output of the monopolist, if she has constant marginal cost of Rs 5. How will the total surplus in this market will change if she practices perfect price discrimination ? 


1
Expert's answer
2022-03-09T11:02:54-0500

"Q=14-2P"


Therefore,


"MC=14-2P"


"MR=14+2P"


"Profit=MR=MC"


But, MC=5


Hence;


"5=14-2P"


"2P=9"


"P=4.5"


The output will be calculated as;


"Q=14-2(4.5)"


"=14-9"


"=9"


If she practices perfect price discrimination buyers will not enjoy consumer surplus at all. The total surplus will consist entirely of producer surplus for the monopoly.